# Berkovich Motives

### v1

@@ 4. Finitary arc-sheaves
is an isomorphism.
@@
Part (i) is clear from the given condition. For part (ii), we rewrite it as the condition that for $A=\mathrm{colim}_{i}A_{i}$ with some initial index $i_{0}$, and a section $x_{0}\in X(A_{i_{0}})$ (inducing compatible sections in $X(A_{i})$ and $X(A)$), one has
$\mathrm{colim}_{i}Y(A_{i})\times_{X(A_{i})}\{x_{0}\}\to Y(A)\times_{X(A)}\{x_{0}\}$
is an isomorphism. This condition is stable under base change. Part (iii) reduces to showing that the diagonal of a finitary map is finitary, which follows from finite limits commuting with filtered colimits. Part (iv) is clear as colimits commute with colimits, and pullback. Given the previous results, it suffices to check (v) for the open immersion $U\subset\mathcal{M}(B)$ given by asking a condition $|T|<r$ for some $T\in B$. This follows from the norm on a filtered colimit being given by an infimum. ∎
@@
By descent, we can assume that $X$ is strictly totally disconnected. In that case $U$ is an increasing union of open and closed subsets $U_{k}\subset U$, with complementary open and closed $Z_{k}\supset Z$. For each $k$, we get an obvious short exact sequence
@@
Let $X$ be nonarchimedean and consider the projection $f:\mathbb{P}^{1}_{X}\to X$ and the constant sheaf $\mathbb{Q}$. Then $f_{\ast}\mathbb{Q}=\mathbb{Q}$: Indeed, this can be checked fibrewise, reducing us to geometric points. There is no Galois cohomology, so this reduces to the cohomology of $\mathcal{M}(\mathbb{P}^{1}_{C})$ which is a tree. If the projection formula were true, this would mean that for any $M\in\mathcal{D}_{\mathrm{fin}}(X,\mathbb{Q})$, one has $f_{\ast}f^{\ast}M=M$ and in particular $M(\mathbb{P}^{1}_{X})=M(X)$. But the sheaf $\overline{\mathbb{G}}_{m}=\mathbb{G}_{m}/(1+\mathcal{O}_{<1})$ studied in the next section yields a counterexample. It is visibly finitary, and concentrated in degree 00 when evaluated at $C$. But $H^{1}(\mathbb{P}^{1}_{C},\overline{\mathbb{G}}_{m})\neq 0$ because $\mathbb{P}^{1}_{C}$ has a nontrivial line bundle. As we will see in the next section, the sheaf $\overline{\mathbb{G}}_{m}$ used in this counterexample is even ball-invariant; without ball-invariance, there are many more counterexamples.
@@ 5. Effective motives
is an isomorphism. By arc-descent, we can assume that $A$ is strictly totally disconnected. We can then treat both as sheaves on the profinite set $\mathcal{M}(A)$. As both sides are finitary (using Theorem 4.20), the stalks agree with the values on the algebraically closed residue fields, where it holds by assumption. ∎
@@
is stable under all all colimits.
@@
The formula defines an endofunctor $L_{\mathbb{B}}$ of $\mathcal{D}_{\mathrm{fin}}(X,\mathbb{Z})$ with a transformation $\mathrm{id}\to L_{\mathbb{B}}$ that is an equivalence on objects of $D_{\mathrm{mot}}^{\mathrm{eff}}(X,\mathbb{Z})$. To see that it is indeed the left adjoint, it suffices to see that $L_{\mathbb{B}}$ takes values in ball-invariant sheaves. Note that
@@
holds also for any analytic Banach ring $A$ of finite cohomological dimension, as then evaluation at $A$ commutes with all colimits in $\mathcal{D}_{\mathrm{fin}}(-,\mathbb{Z})$.
To see that $L_{\mathbb{B}}(M)$ is ball-invariant, we first observe that the two maps
@@
agree. Indeed, those two maps admit a homotopy constructed from a standard homotopy between the two maps $i_{0},i_{1}:\Delta^{\bullet}_{A}\to\Delta^{\bullet}_{\Delta^{1}_{A}}$ of cosimplicial objects. But $i_{0}^{\ast}=i_{1}^{\ast}$ implies that a sheaf is ball-invariant: Indeed, it implies that for any homotopy $f:\Delta^{1}_{A}\times X\to Y$ of maps $X\to Y$, the maps $f_{0}:X\to Y$ and $f_{1}:X\to Y$ induce the same map on sections of the sheaf, and this applies in particular to $f_{0}:\Delta^{1}_{A}\to\Delta^{1}_{A}$ projecting everything to 00, and $f_{1}=\mathrm{id}:\Delta^{1}_{A}\to\Delta^{1}_{A}$, taking $f$ to be multiplication. ∎
@@
The previous proposition then follows by noting that if $M$ is finitary, so is $L_{\mathbb{B}}^{\mathrm{pre}}(M)$; and $L_{\mathbb{B}}(M)$ is its arc-hypersheafification, which then does not change the value on strictly totally disconnected $A$.
@@
for finitary $M$ and $n\geq 0$; and these are $\mathbb{B}$-contractible (via multiplication on $\mathbb{B}^{d}$).
@@
With torsion coefficients prime to the characteristic, étale sheaves are ball-invariant.
@@
Ball-invariance is a statement after pullback to algebraically closed Banach fields, so we can assume that $X=\mathcal{M}(C)$. The assumption on $M$ is then equivalent to the assertion that $M$ is the constant sheaf on the $\mathbb{Z}/n$-module $M_{0}=M(C)$. As taking cohomology on $C\langle T\rangle_{1}$ commutes with arbitrary direct sums, we can assume $M=\mathbb{Z}/n\mathbb{Z}$. We can now appeal to a theorem of Berkovich [Ber93, Theorems 6.3.9, 6.4.1], or argue directly as follows. We compute the higher direct images towards $\mathcal{M}(C\langle T\rangle_{1})$. We know that this is concentrated in degrees 00 and $1$, and in degree 00 we get the constant sheaf $\mathbb{Z}/n\mathbb{Z}$, whose cohomology in turn is just $\mathbb{Z}/n\mathbb{Z}$ in degree 00, as the space is an infinite tree. We need to understand the sheaf in degree $1$, and show that it has no global sections or cohomology. First, we can understand the fibres of the sheaf: These are given by the Galois cohomology groups of the fields $K(x)$. At points of type (4), this is trivial; at points of type (3), given by $\mathbb{Z}/n\mathbb{Z}$; and at points of type (2), given by the reduction modulo $n$ of the group of degree 00 divisors on $\mathbb{P}^{1}_{k}$, where $k$ is the residue field of $C$. As a sheaf, the description is the following: On any finite tree
@@
(for a finite subset $A\subset\mathcal{O}_{C}$), look at the sheaf which locally associates an element of $\mathbb{Z}/n\mathbb{Z}$ to any branch of the tree, so that at branching points the value at the upper branch is the sum of the values at the branches below; and the value at any branch meeting the boundary is zero. Now take the filtered colimit of the pullbacks of these sheaves, where pullback maps assign the value zero to any new branches. Using Kummer covers (and a choice of primitive $n$-th root of unity in $C$) one constructs a map from this candidate sheaf to the first pushforward, and then checks on stalks that it is an isomorphism. To check vanishing of cohomology, reduce to the sheaves on finite trees, and inductively project them to smaller trees. These projection maps are isomorphisms except at one point in the target, where the preimage is some closed interval $I$. This reduces to the vanishing of all cohomology groups of $j_{!}\mathbb{Z}/n\mathbb{Z}$ on $[0,1]$, for $j:[0,1)\hookrightarrow[0,1]$. ∎
With rational coefficients, the category of ball-invariant sheaves also inherits a $t$-structure (the so-called homotopy $t$-structure). However, this behaves differently from the $t$-structure with torsion coefficients; for example the Tate twist will sit in degree $-1$ and not in degree 00.
@@
$0\to H^{1}(C\langle T\rangle_{1},\mathcal{H}^{i-1}M)\to H^{i}(C\langle T\rangle_{1},M)\to H^{0}(C\langle T\rangle_{1},\mathcal{H}^{i}M)\to 0$
@@ 5.1. Free motivic sheaves
We can assume that $X=\mathcal{M}_{\mathrm{arc}}(A)$ is affine. Assume first that $X$ is nonarchimedean. Then the open balls $\mathbb{D}^{n}_{X}(0,<r)\subset\mathbb{A}^{n}_{X}$ around 00 of radius $r$ are subgroups (with respect to coordinatewise addition). Moreover, $U$ is an increasing union of open subspaces $U_{i}$ that are invariant under $\mathbb{D}^{n}_{X}(0,<r_{i})$ for some $r_{i}>0$. Then $U_{i}$ is homotopy-equivalent to the presheaf quotient $U_{i}/\mathbb{D}^{n}_{X}(0,<r_{i})$ (more precisely, to each term in the simplicial presheaf computing it, and hence to its geometric realization), and this quotient is already finitary.
If $X$ is archimedean, then $U$ is a topological manifold and the result is standard. In general, the argument from the nonarchimedean case generalizes at the expense of replacing the actual quotient $U_{i}/\mathbb{D}^{n}_{X}(0,<r_{i})$ by a simplicial approximation. ∎
@@
representing the functor $M\mapsto M(U)$. Moreover, the formation of $\mathbb{Z}_{\mathrm{mot}}[U]$ commutes with base change and sends fibre products to tensor products.
@@
For the final statement, we note that if $\mathcal{D}_{\mathbb{B}}(X_{\mathrm{arc}},\mathbb{Z})\subset\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$ denotes the full subcategory of ball-invariant objects (where ball-invariance means that the values on $Y$ and $Y\times\mathbb{B}$ agree, for all $Y$ over $X$), then the left adjoint to the inclusion defines a symmetric monoidal functor commuting with all pullbacks (as everything is a slice in the arc-site). The previous statement says that when applied to such $U$, this left adjoint takes values in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)\subset\mathcal{D}_{\mathbb{B}}(X_{\mathrm{arc}},\mathbb{Z})$, and this inclusion does commute with tensor products and pullbacks. ∎
@@
Writing $f$ as a composite, we can assume that $Y=\mathbb{A}^{n}_{X}$. We can assume that $X$ is strictly totally disconnected. Then $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathbb{A}^{n}_{X})$ is generated by $\mathbb{Z}_{\mathrm{mot}}[V_{/Y}]$ where $V\subset\mathbb{A}^{m}_{Y}$ is open. But then $f_{\sharp}(\mathbb{Z}_{\mathrm{mot}}[V_{/Y}])=\mathbb{Z}_{\mathrm{mot}}[V_{/X}]$, where $V\subset\mathbb{A}^{n+m}_{X}$ is also open. This description commutes with base change, and satisfies the projection formula (where one has to check the latter only on generators $\mathbb{Z}_{\mathrm{mot}}[U_{/X}]$, where it follows from $\mathbb{Z}_{\mathrm{mot}}$ taking fibre products to tensor products). ∎

### v2

@@ 4. Finitary arc-sheaves
is an isomorphism. Equivalently, the diagram
${\mathrm{colim}_{i}Y(A_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{colim}_{i}X(A_{i})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${X(A)}$
is cartesian.
We note that this is equivalently the condition that $Y$ defines a finitary arc-sheaf over $X$.
@@
Parts (i) – (iv) follow directly from the definition and standard properties of anima. Given these results, it suffices to check (v) for the open immersion $U\subset\mathcal{M}(B)$ given by asking a condition $|T|<r$ for some $T\in B$. This follows from the norm on a filtered colimit being given by an infimum. For (vi), we can assume that $f:Y\to X$ has strictly totally disconnected target. For any $x\in X$ in the image of $f$, we can write $x$ as a cofiltered limit of open and closed neighborhoods, and by finitaryness, one such neighborhood must lie in $Y$. It follows that $Y$ is open. ∎
@@
By descent, we can assume that $X$ is strictly totally disconnected. In that case $U$ is an increasing union of open and closed subsets $U_{k}\subset U$, with complementary open and closed $Z_{k}\supset Z$ (for $k$ in some filtered index category $K$). For each $k$, we get an obvious short exact sequence
@@
Let $X$ be nonarchimedean and consider the projection $f:\mathbb{P}^{1}_{X}\to X$ and the constant sheaf $\mathbb{Q}$. Then $f_{\ast}\mathbb{Q}=\mathbb{Q}$: Indeed, this can be checked fibrewise, reducing us to geometric points. There is no Galois cohomology, so this reduces to the cohomology of $\mathcal{M}(\mathbb{P}^{1}_{C})$ which is a tree. If the projection formula were true, this would mean that for any $M\in\mathcal{D}_{\mathrm{fin}}(X,\mathbb{Q})$, one has $f_{\ast}f^{\ast}M=M$ and in particular $M(\mathbb{P}^{1}_{X})=M(X)$. But the sheaf $\overline{\mathbb{G}}_{m}=\mathbb{G}_{m}/(1+\mathcal{O}_{<1})$ studied in the next section yields a counterexample. It is visibly finitary, and concentrated in degree 0 when evaluated at $C$. But $H^{1}(\mathbb{P}^{1}_{C},\overline{\mathbb{G}}_{m})\neq 0$ because $\mathbb{P}^{1}_{C}$ has a nontrivial line bundle. As we will see in the next section, the sheaf $\overline{\mathbb{G}}_{m}$ used in this counterexample is even ball-invariant; without ball-invariance, there are many more counterexamples.
@@ 5. Effective motives
is an isomorphism. By arc-descent, we can assume that $A$ is strictly totally disconnected. We can then treat both sides as sheaves on the profinite set $\mathcal{M}(A)$. As both sides are finitary (using Theorem 4.20), the stalks agree with the values on the algebraically closed residue fields, where the result holds by assumption. ∎
@@
is stable under all colimits.
@@
Using Theorem 4.20, the formula defines an endofunctor $L_{\mathbb{B}}$ of $\mathcal{D}_{\mathrm{fin}}(X,\mathbb{Z})$ with a transformation $\mathrm{id}\to L_{\mathbb{B}}$ that is an equivalence on objects of $D_{\mathrm{mot}}^{\mathrm{eff}}(X,\mathbb{Z})$. To see that it is indeed the left adjoint, it suffices to see that $L_{\mathbb{B}}$ takes values in ball-invariant sheaves, and that the natural transformation $L_{\mathbb{B}}\to L_{\mathbb{B}}^{2}$ obtained by applying $L_{\mathbb{B}}$ to $\mathrm{id}\to L_{\mathbb{B}}$ is an equivalence; see [Lur09, Proposition 5.2.7.4].
Note that
@@
holds also for any analytic Banach ring $A$ of finite cohomological dimension, as then evaluation at $A$ commutes with all colimits in $\mathcal{D}_{\mathrm{fin}}(-,\mathbb{Z})$. To see that $L_{\mathbb{B}}(M)$ is ball-invariant, we first observe that the two maps
@@
agree. Indeed, those two maps admit a homotopy constructed from a standard homotopy between the two maps $i_{0},i_{1}:\Delta^{\bullet}_{A}\to\Delta^{\bullet}_{\Delta^{1}_{A}}$ of cosimplicial objects. But $i_{0}^{\ast}=i_{1}^{\ast}$ implies that a sheaf is ball-invariant: Indeed, it implies that for any homotopy $f:\Delta^{1}_{A}\times X\to Y$ of maps $X\to Y$, the maps $f_{0}:X\to Y$ and $f_{1}:X\to Y$ induce the same map on sections of the sheaf, and this applies in particular to $f_{0}:\Delta^{1}_{A}\to\Delta^{1}_{A}$ projecting everything to 0, and $f_{1}=\mathrm{id}:\Delta^{1}_{A}\to\Delta^{1}_{A}$, taking $f$ to be multiplication.
Finally, the map $L_{\mathbb{B}}\to L_{\mathbb{B}}^{2}$ is an equivalence by another homotopy between cosimplicial objects. ∎
@@
The previous proposition then follows by noting that if $M$ is a finitary arc-presheaf, so is $L_{\mathbb{B}}^{\mathrm{pre}}(M)$; and $L_{\mathbb{B}}(M)$ is its arc-hypersheafification, which then does not change the value on strictly totally disconnected $A$.
@@
In the following proposition, we say that $M$ is $\mathbb{B}$-contractible if there is a map $\mathbb{B}\to\mathrm{End}(M)$ sending $1$ to $1$ and 0 to 0.
@@
for finitary $M$ and $n\geq 0$; and these are $\mathbb{B}$-contractible (via multiplication on $\mathbb{B}^{n}$).
@@
With torsion coefficients prime to the characteristic, étale sheaves are ball-invariant. We did not define the étale site here, and instead use the condition “invariance under change of algebraically closed field” to single out étale sheaves inside all finitary arc-sheaves.
@@
Ball-invariance is a statement after pullback to algebraically closed Banach fields, so we can assume that $X=\mathcal{M}(C)$. The assumption on $M$ is then equivalent to the assertion that $M$ is the constant sheaf on the $\mathbb{Z}/n$-module $M_{0}=M(C)$. As taking cohomology on $C\langle T\rangle_{1}$ commutes with arbitrary direct sums, we can assume $M=\mathbb{Z}/n\mathbb{Z}$. We can now appeal to a theorem of Berkovich [Ber93, Theorems 6.3.9, 6.4.1], or argue directly as follows. We compute the higher direct images towards $\mathcal{M}(C\langle T\rangle_{1})$. We know that this is concentrated in degrees 0 and $1$, and in degree 0 we get the constant sheaf $\mathbb{Z}/n\mathbb{Z}$, whose cohomology in turn is just $\mathbb{Z}/n\mathbb{Z}$ in degree 0, as the space is an infinite tree. We need to understand the sheaf in degree $1$, and show that it has no global sections or cohomology. First, we can understand the fibres of the sheaf: These are given by the Galois cohomology groups of the fields $K(x)$. At points of type (1) or (4), this is trivial; at points of type (3), given by $\mathbb{Z}/n\mathbb{Z}$; and at points of type (2), given by the reduction modulo $n$ of the group of degree 0 divisors on $\mathbb{P}^{1}_{k}$, where $k$ is the residue field of $C$. As a sheaf, the description is the following: On any finite tree
@@
(for a finite subset $A\subset\mathcal{O}_{C}$), look at the sheaf which locally associates an element of $\mathbb{Z}/n\mathbb{Z}$ to any branch of the tree, so that at branching points the value at the upper branch is the sum of the values at the branches below; and the value at any branch meeting the boundary is zero. Now take the filtered colimit of the pullbacks of these sheaves, where pullback maps assign the value zero to any new branches. Using Kummer covers (and a choice of primitive $n$-th root of unity in $C$) one constructs a map from this candidate sheaf to the first pushforward, and then checks on stalks that it is an isomorphism. For the construction of the map, pick any such finite tree, given by some subset $A$ of $\mathcal{O}_{C}$. On any branch of the tree, the sheaf is the constant sheaf $\mathbb{Z}/n\mathbb{Z}$, and we send a class $k\in\mathbb{Z}/n\mathbb{Z}$ to the Kummer cover given by $n$-th roots of the function $(T-a)^{k}$ (where $a$ is chosen so that the branch of the tree is centered at $a$). At a branch point of the tree, we have the branches below, centered at $a_{i}$, and with assigned values $k_{i}\in\mathbb{Z}/n\mathbb{Z}$. We take this to the Kummer cover given by $\prod_{i}(T-a_{i})^{k_{i}}$. On all the branches below, this gives the correct Kummer cover already defined, as when $T$ is close to $a_{j}$ for $j\neq i$ then
$\frac{T-a_{i}}{a_{j}-a_{i}}=1+\frac{T-a_{j}}{a_{j}-a_{i}}$
is close to $1$ and hence an $n$-th power; while $a_{j}-a_{i}\in C$ is also an $n$-th power, so the Kummer cover defined by $T-a_{i}$ becomes trivial. Thus, the Kummer cover for $\prod_{i}(T-a_{i})^{k_{i}}$ agrees with the Kummer cover for $(T-a_{j})^{k_{j}}$ when $T$ is close to $a_{j}$.
Also, at the branch above the branch point, the Kummer cover for $\prod_{i}(T-a_{i})^{k_{i}}$ becomes equivalent to the one for $(T-a)^{\sum k_{i}}$ where $a$ is given by any of the $a_{i}$. Indeed,
$\frac{T-a_{i}}{T-a}=1+\frac{a-a_{i}}{T-a_{i}}$
is close to $1$ in this case, and hence an $n$-th power. This computation reflects the condition that the value at the upper branch must be the sum of the values at the lower branches.
This constructs the map from this candidate sheaf towards the first pushforward, and one directly checks that at stalks, one gets an isomorphism, yielding thus an explicit description of this first pushforward.
To check vanishing of cohomology of this sheaf, reduce (by filtered colimits) to the sheaves on finite trees, and inductively project them to smaller trees. These projection maps are isomorphisms except at one point in the target, where the preimage is some closed interval $I$. This reduces to the vanishing of all cohomology groups of $j_{!}\mathbb{Z}/n\mathbb{Z}$ on $[0,1]$, for $j:[0,1)\hookrightarrow[0,1]$. ∎
With rational coefficients, the category of ball-invariant sheaves also inherits a $t$-structure (the so-called homotopy $t$-structure). However, this behaves differently from the $t$-structure with torsion coefficients; for example the Tate twist will sit in degree 1-1 and not in degree 0.
@@
$0\to H^{1}(C\langle T\rangle_{1},\mathcal{H}^{i-1}M)\to H^{i}(C\langle T\rangle_{1},M)\to H^{0}(C\langle T\rangle_{1},\mathcal{H}^{i}M)\to 0.$
@@ 5.1. Free motivic sheaves
We can assume that $X=\mathcal{M}_{\mathrm{arc}}(A)$ is affine. Assume first that $X$ is nonarchimedean. Then the open balls $\mathbb{D}^{n}_{X}(0,<r)\subset\mathbb{A}^{n}_{X}$ around 0 of radius $r$ are subgroups (with respect to coordinatewise addition). Moreover, $U$ is an increasing union of open subspaces $U_{i}$ that are invariant under $\mathbb{D}^{n}_{X}(0,<r_{i})$ for some $r_{i}>0$. Then $U_{i}$ is homotopy-equivalent to the presheaf quotient $U_{i}/\mathbb{D}^{n}_{X}(0,<r_{i})$ (more precisely, to each term in the simplicial presheaf computing it, and hence to its geometric realization), and this quotient is already finitary.
In general, we argue as follows to prove finitarity. Consider any filtered colimit of Banach rings $A=\mathrm{colim}_{i}A_{i}$ over $X$. We need to see that the map of simplicial sets
$\mathrm{colim}_{i}U(A_{i}\langle\Delta^{\bullet}\rangle)\to U(A\langle\Delta^{\bullet}\rangle)$
induces an equivalence of anima, under geometric realization. Using Kan’s $\mathrm{Ex}^{\infty}$-functor, this can be checked by mapping simplicial subdivisions $\mathrm{sd}^{\ell}\partial\Delta^{n}$ of $\partial\Delta^{n}$’s into these simplicial sets, and checking that it induces homotopy equivalences. The key now is that any two maps that are sufficiently close are automatically $\mathbb{B}$-homotopic, via the straight-line interpolation in $U\subset\mathbb{A}^{n}$.
More precisely, given any finite simplicial set $S$ with a map $f:S\to U(A\langle\Delta^{\bullet}\rangle)$ as well as a sub-simplicial set $S^{\prime}\subset S$ with a sequence of maps $g^{\prime}_{n}:S^{\prime}\to U(A_{i_{n}}\langle\Delta^{\bullet}\rangle)$ such that the $g^{\prime}_{n}$ converge to $f|_{S^{\prime}}$, we can always find some $j_{n}\geq i_{n}$ and an extension of $g^{\prime}_{n}$ to a map $g_{n}:S\to U(A_{j_{n}}\langle\Delta^{\bullet}\rangle)$ such that the $g_{n}$ converge to $f$. Indeed, this can as usual be done by induction on simplices, reducing to $S=\Delta^{k}$ and $S^{\prime}=\partial\Delta^{k}$, where it follows from the map
$\mathrm{colim}_{i}\mathrm{ker}(A_{i}\langle\Delta^{k}\rangle\to A_{i}\langle\partial\Delta^{k}\rangle)\to\mathrm{ker}(A\langle\Delta^{k}\rangle\to A\langle\partial\Delta^{k}\rangle)$
having dense image (and $U$ being open in affine space). Thus, given any map $f:\mathrm{sd}^{\ell}\partial\Delta^{k}\to U(A\langle\Delta^{\bullet}\rangle)$, we can find a sequence $g_{n}$ of maps $\mathrm{sd}^{\ell}\partial\Delta^{k}\to U(A_{i_{n}}\langle\Delta^{\bullet}\rangle)$ so that the $g_{n}$ converge to $f$. In particular, for sufficiently large $n$, the straight-line interpolation from $f$ to $g_{n}$ defines a $\mathbb{B}$-homotopy in maps from $\mathrm{sd}^{\ell}\partial\Delta^{k}$ to $U(A\langle\Delta^{\bullet}\rangle)$, showing surjectivity of
$\mathrm{colim}_{i}U(A_{i}\langle\Delta^{\bullet}\rangle)\to U(A\langle\Delta^{\bullet}\rangle)$
on homotopy groups. For injectivity on homotopy groups, we argue similarly, using maps from $\mathrm{sd}^{\ell}(\Delta^{1}\times\partial\Delta^{k})$ with fixed restriction to $\mathrm{sd}^{\ell}(\{0,1\}\times\partial\Delta^{k})$. ∎
@@
representing the functor $M\mapsto M(U)$. Moreover, the formation of $\mathbb{Z}_{\mathrm{mot}}[U]$ commutes with base change and sends fibre products over $X$ to tensor products.
@@
For the final statement, we note that if $\mathcal{D}_{\mathbb{B}}(X_{\mathrm{arc}},\mathbb{Z})\subset\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$ denotes the full subcategory of ball-invariant objects (where ball-invariance means that the values on $Y$ and $Y\times\mathbb{B}$ agree, for all $Y$ over $X$), then the left adjoint to the inclusion defines a symmetric monoidal functor commuting with all pullbacks. Indeed, $\mathcal{D}_{\mathbb{B}}(X_{\mathrm{arc}},\mathbb{Z})$ is the symmetric monoidal Verdier quotient of $\mathcal{D}(X_{\mathrm{arc}},\mathbb{Z})$ by the $\otimes$-ideal generated by the cone of $\mathbb{Z}[\mathbb{B}]\to Z\mathbb{Z}$; and the commutation with pullbacks is left adjoint to the observation that pushforward preserves ball-invariance.
Now the previous lemma says that when applied to $U$ as in the statement, this left adjoint takes values in $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(X)\subset\mathcal{D}_{\mathbb{B}}(X_{\mathrm{arc}},\mathbb{Z})$. This subcategory is stable under tensor products and pullbacks, giving the desired statement. ∎
@@
Writing $f$ as a composite and using Proposition 4.24, we can assume that $Y=\mathbb{A}^{n}_{X}$. We can assume that $X$ is strictly totally disconnected; indeed, if we can prove the left adjoint exists in that case, and commutes with any base change of strictly totally disconnected spaces, then we get existence in general by hyperdescent (using [Lur17, Proposition 4.7.4.18]). If $X$ is strictly totally disconnected, $\mathcal{D}_{\mathrm{mot}}^{\mathrm{eff}}(\mathbb{A}^{n}_{X})$ is generated by $\mathbb{Z}_{\mathrm{mot}}[V_{/Y}]$ where $V\subset\mathbb{A}^{m}_{Y}$ is open. But then $f_{\sharp}(\mathbb{Z}_{\mathrm{mot}}[V_{/Y}])=\mathbb{Z}_{\mathrm{mot}}[V_{/X}]$, where $V\subset\mathbb{A}^{n+m}_{X}$ is also open. This description commutes with base change, and satisfies the projection formula (where one has to check the latter only on generators $\mathbb{Z}_{\mathrm{mot}}[U_{/X}]$, where it follows from $\mathbb{Z}_{\mathrm{mot}}$ taking fibre products over $XC$ to tensor products). ∎